OpenStudy (anonymous):

If a = 2 and b = -10, what is the value of the expression below? 5a + b/a + 2a

OpenStudy (anonymous):

A 9 B 10 C 6

OpenStudy (anonymous):

C

OpenStudy (anonymous):

is this a test question or homework

OpenStudy (anonymous):

but actually -6

OpenStudy (raden):

5a + b/a + 2a = 5(2) + (-10/2) + 2(2) = 10 - 5 + 4 = 9